# Qingchuan Yang 1,  Simon Mahns  [^2] footnotemark: 2,  Sida Li 322footnotemark: 2,  Anri Gu 322footnotemark: 2,  Jibang Wu 4,  Haifeng Xu  [^3] footnotemark: 3

### v1

@@ 3.1.3 Market Return to Measure Economic Values
Prophet Arena  uses events from real-world prediction markets, hence allows us to evaluate the economic value of a forecast based on current market prices. We thus introduce Average Return as a natural metric to capture how profitable it would be to trade in the market using LLM forecasts? Formally, consider a (binary) market with market price $q^{Y}$ per share for the Yes contract and $q^{N}$ for the No contract. Given forecasted probability $p$ for an Yes outcome, we use the following natural betting strategy: allocate a unit budget (\$1) to buy $1/q_{Y}$ shares of Yes contracts if $\frac{p}{q^{Y}}\geq\frac{1-p}{q^{N}}$, or to buy $1/q^{N}$ shares of No contracts otherwise. After the event resolution, if the bought contracts match the event outcome, the return is simply the number of these contracts; otherwise, the return is 0. The Average Return of an LLM is then defined as the average of these returns across all markets, under a unit budget allocated to each market.
Remark: Three notes on market return. First, the betting strategy above can be shown to maximize expected return under risk neutrality. We defer its proof to § B.5, where we describe the more general optimization framework for betting strategies under different risk-preferences. Second, while both market return and Brier score assess forecast quality, they differ fundamentally. The Brier score is an absolute metric, measuring a forecast’s closeness to the ground truth, independent of external factors like market prices. In contrast, market return is a relative metric, capturing how much a forecast outperforms the market’s current belief (interpreted as the contract price (Wolfers & Zitzewitz, 2006)). This distinction is already evident in our evaluations. In § B.4, we provide a concrete example to illustrate that forecasts with worse Brier scores can achieve higher market returns. Third, calibration is not related to the total market return, but is indicative about how balanced the market return is from betting on the Yes contracts and No contracts. To formalize this intuition, we prove in § B.6 that a well-calibrated and symmetric forecaster – intuitively, one that is not systematically more aggressive or conservative than the market – will have balanced returns from both contract types.
@@ 3.2.2 Evaluation Results and Discussions
Overall, our results suggest that absolute forecasting skill and relative profitability against prediction markets are still challenging for today’s LLMs. A more fine-grained investigation could help deepen our understanding about what could make an LLM a good prophet. Fig. 3 draws for the best and worst models (Left & right) their reliability diagrams regarding calibration errors (Guo et al., 2025), where predicted probability ($x$-axis) is compared against realized frequency ($y$-axis) at different probability bins. While their calibration is similar in intermediate ranges, the stronger model – OpenAI o3 – performs much better in the extreme bins (0-$0.1$ and 0.90.9-$1.0$), where it almost always predicts correctly. Because such extreme forecasts occur frequently, this advantage helps explain the gap in both Brier score and market return.
@@ 4.2.1 Can Internalized Knowledge Become Foresight?
In Fig. 4, we observe that models most reliably recall events in Entertainment. By contrast, Climate and Weather and Politics display low recall and frequent mis-recall. Two factors likely contribute. First, Weather prompts often require fine-grained, date-stamped facts (e.g., “Highest temperature in Miami on Aug 29, 2023?”), which are less likely to be memorized. Second, due to 2023 regulatory constraints on Kalshi’s election markets, Politics in our dataset skews toward politics-adjacent indicators (e.g., “Biden 538 approval rating on Aug 30, 2023?”), which varies daily and may be sparsely represented in training corpora (Inc. (2025)).
@@ 4.3 Granular Analysis of LLM-as-a-Prophet
The reported probabilities of LLMs compress a rich decision-making process into a single number: two models may yield similar predicted probabilities and scores while relying on dramatically different reasoning processes. To address this, we open the black box and evaluate the process underlying each forecast. Specifically, we adopt an LLM-as-a-judge framework (Zheng et al., 2023) to assess the soundness of reasoning along five critical dimensions: source selection, evidence extraction, reasoning synthesis, reasoning-to-prediction alignment, and recognition of prediction uncertainty. Full prompts and additional details of the evaluation framework are provided in § C.6.
@@ A.5 Comprehensive List of Evaluated LLMs
GPT-5 OpenAI (2025b) No High
o3 OpenAI (2025c) No High
o3-Mini OpenAI (2025c) No Medium
o4-Mini OpenAI (2025c) No High
Gemini 2.5 Pro Comanici et al. (2025) No Enabled
Gemini 2.5 Flash Comanici et al. (2025) No Enabled
Grok-4 xAI (2025b) No Enabled
Grok-3-Mini xAI (2025a) No Enabled
GPT-4.1 OpenAI (2025a) No N/A
Claude Sonnet 4 Anthropic (2025) No Enabled
Kimi-K2 Kimi Team (2025) Yes N/A
GPT-4o OpenAI (2024) No N/A
Llama 4 Maverick Meta Team (2025) Yes N/A
Llama 4 Scout Meta Team (2025) Yes N/A
DeepSeek-V3 Liu et al. (2024) Yes N/A
DeepSeek-R1 Guo et al. (2025) Yes Enabled
Qwen3-235B Yang et al. (2025) Yes N/A
Gemini 2.0 Flash Gemini Team (2024) No N/A
Gemini 2.0 Flash (Lite) Gemini Team (2024) No N/A
@@ B.1 Further Intuitions behind Brier Score
B.2 A Better-calibrated forecast can have a worse Brier score.
As mentioned in § 3.1.2, calibration errors and Brier scores measure different aspects of forecasts. This difference can be illustrated with a simple example. Suppose there are two markets $M_{1},M_{2}$ with ground-truth probability $p^{*}_{1}=0.9,p^{*}_{2}=0.1$. Alice’s prediction is $p^{A}_{1}=1,p^{A}_{2}=0$ whereas Bob’s prediction is $p^{B}=0.5$ for both markets. It is not difficult to see that Alice has a better/smaller Brier score, but she turns out to have a worse/larger ECE. Specifically, since Alice has two different probability predictions, her ECE is $\frac{1}{2}(|p^{*}_{1}-p^{A}_{1}|+|p^{*}_{2}-p^{A}_{2}|)=0.1$ whereas Bob has a single probability prediction with the ECE equaling $\frac{1}{2}\big|p^{*}_{1}+p^{*}_{2}-2p^{B}|\big|=0$. That is, Bob’s prediction of $0.5$ probability is perfectly calibrated and trustworthy as, conditioned on this probability, events’ average probability is indeed $0.5$.
B.3 Empirical Estimation of Calibration Error
@@ B.5 A Unified Framework for Utility-Maximizing Betting Strategy
Contracts and market prices. Each binary market resolves to either Yes or No. A share of a Yes contract pays \$1 if the outcome is Yes and \$0 otherwise; the price of this contract is denoted $q$. Symmetrically, a share of a No contract costs 1q1-q and pays \$1 if the outcome is No. These prices are often interpreted as market-implied probabilities (Wolfers & Zitzewitz, 2006), though our framework does not rely on this interpretation.
@@ B.7 Bootstrap Confidence Intervals
To assess the uncertainty of the resulting point estimate, we construct a nonparametric bootstrap confidence interval (CI) at level $(1-\alpha)\%$ (DiCiccio & Efron, 1996). Specifically, we form a symmetric interval around the point estimate using bootstrap resampling. A pseudo-code for our implementation is given below in Algorithm 1.
@@ C.2 Sharpe Ratio for Average Return
As discussed in § 3.2 and Table 5, Average Return often exhibits wide confidence intervals, reflecting the high variance of event-level payoffs. In most events, models earn nothing or only marginal gains, while in rare cases they achieve outsized returns. To account for this variability, we complement Average Return with the Sharpe ratio (Sharpe, 1998), a standard metric that normalizes expected returns by their volatility. Formally, for asset returns $R_{a}$, the Sharpe ratio is defined as
@@ C.3 Probability Elicitation Methods
    Verbalized Probability (Tian et al., 2023)
@@
    • Prompt emsemble (Wightman et al., 2023): for each market, we average the probabilities elicited from the default and all variation prompts (i.e. (A), (B), (C) above).
@@
    Self-consistency (Wang et al., 2022)
@@
    • Weighted (Taubenfeld et al., 2025): we further supplement each Yes/No with a confidence score. [^11] This confidence reflects uncertainty about the chosen answer, not a direct market probability (e.g., low confidence in Yes signals indecision, not belief in “NO”). Probabilities are formed by confidence-weighted aggregation.
@@ C.6 Evaluating Reasoning using an LLM-as-a-Judge Framework
To directly evaluate LLMs reasoning processes, we employ an LLM-as-a-judge framework (Zheng et al., 2023) to assess the soundness of their reasoning methodologies. Our evaluation proceeds in two stages: (1) eliciting explicit reasoning from the prediction model, and (2) systematically evaluating this reasoning with an independent evaluator.
@@
C.6.1 (Lack of) Effect of Scaffolding Prompt on Prediction

### v2

@@ 3.1.3 Market Return to Measure Economic Values
Prophet Arena  uses events from real-world prediction markets, hence allows us to evaluate the economic value of a forecast based on current market prices. We thus introduce Average Return as a natural metric to capture how profitable it would be to trade in the market using LLM forecasts? Formally, consider a (binary) market with market price $q^{Y}$ per share for the Yes contract and $q^{N}$ for the No contract. Given forecasted probability $p$ for an Yes outcome, we use the following natural betting strategy: allocate a unit budget (\$1) to buy $1/q_{Y}$ shares of Yes contracts if $\frac{p}{q^{Y}}\geq\frac{1-p}{q^{N}}$, or to buy $1/q^{N}$ shares of No contracts otherwise. [^7] This betting strategy can be shown to maximize expected return under risk neutrality. We defer its proof to § B.5, describing a general framework for optimizing betting strategies under different risk-preferences. After the event resolution, if the bought contracts match the event outcome, the return is simply the number of these contracts; otherwise, the return is 0. The Average Return of an LLM is then defined as the average of these returns across all markets, under a unit budget allocated to each market.
Remark: Intrinsic Differences among the Three Metrics. First, the difference between the Brier score and ECE is well-known (murphy1973new).888murphy1973new shows that the Brier score can be decomposed into three terms: the ECE (also coined it “reliability score” by murphy1973new), the uncertainty score that captures inherent randomness of the to-be-forecasted event, and the resolution score that captures forecasts’ variance. A forecast with better/smaller ECE could have worse Brier score (see an example in § B.3). In practice, ECE is crucial when decision makers have risk preferences since smaller ECE ensures that the risk encoded in the forecasted probabilities is more reliably captured, leading to better risk-preference-adjusted decisions – even when the forecasted probabilities have worse Brier scores. In § B.3, we illustrate this insight with a simple example in which a forecaster with better ECE but much worse Brier score could lead to better risk-adjusted utility. Second, while both market return and Brier score assess forecast quality, they differ fundamentally. The Brier score is an absolute metric, measuring a forecast’s closeness to the ground truth, independent of external factors like market prices. In contrast, market return is a relative metric, capturing how much a forecast outperforms the market’s current belief (interpreted as the contract price (wolfers2006interpreting)). This distinction is already evident in our evaluations. In § B.4, we provide a concrete example to illustrate that forecasts with worse Brier scores can achieve higher market returns. Third, calibration is not related to the total market return, but is indicative about how balanced the market return is from betting on the Yes contracts and No contracts. To formalize this intuition, we prove in § B.6 that a well-calibrated and symmetric forecaster – intuitively, one that is not systematically more aggressive or conservative than the market – will have balanced returns from both contract types.
@@ 3.2.2 Evaluation Results and Discussions
Overall, our results suggest that absolute forecasting skill and relative profitability against prediction markets are still challenging for today’s LLMs. A more fine-grained investigation could help deepen our understanding about what could make an LLM a good prophet. Fig. 3 draws for the best and worst models (Left & right) their reliability diagrams regarding calibration errors (guo2025deepseek), where predicted probability ($x$-axis) is compared against realized frequency ($y$-axis) at different probability bins. While their calibration is similar in intermediate ranges, the stronger model – OpenAI o3 – performs much better in the extreme bins (0-$0.1$ and 0.90.9-$1.0$), where it almost always predicts correctly. Because such extreme forecasts occur frequently, this advantage helps explain the gap in both Brier score and market return.
The empirical results reported in this paper are based on evaluations conducted over 1,367 events resolved prior to October 11, 2025. These events were collected via our real-time Kalshi data pipeline, and consequently the distribution of event categories reflects the underlying composition: 76% Sports, 8% Entertainment, 7% Politics, and 9% Other. To ensure that the results remain robust to event composition, we additionally verify that these performance patterns are robust to reweighting the dataset toward a more balanced distribution across categories. Full results from these balanced-subset evaluations are reported in Appendix B.8. Regardless, we plan to continue expanding and diversifying the event pool in future releases of Prophet Arena.
@@ 4.2.1 Can Internalized Knowledge Become Foresight?
In Fig. 4, we observe that models most reliably recall events in Entertainment. By contrast, Climate and Weather and Politics display low recall and frequent mis-recall. Two factors likely contribute. First, Weather prompts often require fine-grained, date-stamped facts (e.g., “Highest temperature in Miami on Aug 29, 2023?”), which are less likely to be memorized. Second, due to 2023 regulatory constraints on Kalshi’s election markets, Politics in our dataset skews toward politics-adjacent indicators (e.g., “Biden 538 approval rating on Aug 30, 2023?”), which varies daily and may be sparsely represented in training corpora (kalshi).
@@ 4.3 Granular Analysis of LLM-as-a-Prophet
The reported probabilities of LLMs compress a rich decision-making process into a single number: two models may yield similar predicted probabilities and scores while relying on dramatically different reasoning processes. To address this, we open the black box and evaluate the process underlying each forecast. Specifically, we adopt an LLM-as-a-judge framework (zheng2023judging) to assess the soundness of reasoning along five critical dimensions: source selection, evidence extraction, reasoning synthesis, reasoning-to-prediction alignment, and recognition of prediction uncertainty. Full prompts and additional details of the evaluation framework are provided in § C.6.
To assess the reliability of the LLM-as-a-judge framework, we conducted a blinded human evaluation on a random subset of 170 predictions made by the LLMs. Human raters scored the same five dimensions using the identical rubric provided to the LLM judge. Overall, human and LLM ratings were closely aligned (mean difference <0.5<0.5 out of a maximum of 4, with standard deviation $\approx 0.3$ across all dimensions). Full details of the study design and results are provided in § C.6.1.
@@ 5 Conclusion
$\clubsuit$ Appendix: Table of Contents
  1. A.
    Prophet Arena Pipeline Details ........................................................................................................................................................................Page A
  2. B.
    Technical Details ........................................................................................................................................................................Page B
  3. C.
    Additional Experiment Results ........................................................................................................................................................................Page C
  4. D.
    Case Studies ........................................................................................................................................................................Page D
  5. E.
    Prompts ........................................................................................................................................................................Page E
@@ A.5 Comprehensive List of Evaluated LLMs
GPT-5 gpt5 No High
o3 o3 No High
o3-Mini o3 No Medium
o4-Mini o3 No High
Gemini 2.5 Pro comanici2025gemini No Enabled
Gemini 2.5 Flash comanici2025gemini No Enabled
Grok-4 xai-grok4 No Enabled
Grok-3-Mini xai-grok3 No Enabled
GPT-4.1 gpt41 No N/A
Claude Sonnet 4 claude4 No Enabled
Kimi-K2 team2025kimi Yes N/A
GPT-4o gpt4o No N/A
Llama 4 Maverick meta2025llama4 Yes N/A
Llama 4 Scout meta2025llama4 Yes N/A
DeepSeek-V3 liu2024deepseek Yes N/A
DeepSeek-R1 guo2025deepseek Yes Enabled
Qwen3-235B yang2025qwen3 Yes N/A
Gemini 2.0 Flash gemini20 No N/A
Gemini 2.0 Flash (Lite) gemini20 No N/A
@@ B.1 Further Intuitions behind Brier Score
B.2 Empirical Estimation of Calibration Error
@@
B.3 Brier Score vs ECE.
As discussed in § 3.1, calibration errors and Brier scores measure different aspects of forecasts. We start with a simple example to illustrate that a better-calibrated forecast can have a worse Brier score. Suppose there are two markets $M_{1},M_{2}$ with ground-truth probability $p^{*}_{1}=0.9,p^{*}_{2}=0.1$. Alice’s prediction is $p^{A}_{1}=1,p^{A}_{2}=0$ whereas Bob’s prediction is $p^{B}=0.5$ for both markets. It is not difficult to see that Alice has a better/smaller Brier score, but she turns out to have a worse/larger ECE. Specifically, since Alice has two different probability predictions, her ECE is $\frac{1}{2}(|p^{*}_{1}-p^{A}_{1}|+|p^{*}_{2}-p^{A}_{2}|)=0.1$ whereas Bob has a single probability prediction with the ECE equaling $\frac{1}{2}\big|p^{*}_{1}+p^{*}_{2}-2p^{B}|\big|=0$. That is, Bob’s prediction of $0.5$ probability is perfectly calibrated and reliable as, conditioned on this probability, events’ average probability is indeed $0.5$.
A forecaster with smaller ECE but larger Brier score induces better risk-adjusted decision making. Consider an event $E$ with binary outcome $o\in\{0,1\}$ and a ground-truth probability $p^{*}=0.5$ to have $o=1$. Suppose forecaster $A$ predicts probability $p_{A}:=\Pr(o=1)$ which correlates with the realized outcome $o$ in the following way
[…]
where $\epsilon$ is a small value. In other words, forecaster $A$ predicts $p_{A}=1$ or $p_{A}=0$, each with probability $0.5$ and with $\epsilon$ fraction of errors. These are extremal and uncalibrated forecasts but is almost always correct. It is easy to verify that the Brier score of $p_{A}$ is $0.5\epsilon\times(1-0)^{2}+0.5\epsilon\times(1-0)^{2}=\epsilon$, whereas expected calibration error (ECE) is $0.5\times[1-(1-\epsilon)]^{2}+0.5\times[1-(1-\epsilon)]^{2}=0.5\epsilon^{2}$.
Next consider forecaster $B$ predicts probability $p_{B}:=\Pr(o=1)$ which correlates with the realized outcome $o$ in the following way
[…]
That is, forecaster $B$ predicts $p_{B}=2/3$ for $75\%$ of the time, and predicts $p_{B}=0$ otherwise. It is easy to see that $p_{B}$ is perfectly calibrated with $ECE=0$, but has Brier score equal $0.75\times[\frac{2}{3}(2/3-1)^{2}+\frac{1}{3}(2/3-0)^{2}]=1/6$.
Clearly, $p_{B}$ is more well-calibrated than $p_{A}$ but has much worse Brier score, especially when $\epsilon$ is small. However, we show that $p_{B}$ is better for risk-averse decision making, than $p_{A}$. Suppose a decision maker has $1$ unit of fund and would like to buy contracts for event $E$’s outcome. Suppose the current market price for the contract $o=1$ and $o=0$ are both $0.5$ per contract. Adopting a classic risk-averse utility function, we assume her utility $U(x)=\log(x)$ (i.e., $x$ amount of return means $U(x)$ utility to her).
If the decision maker adopts forecaster $A$, since $A$ makes extremal forecasts, even under risk-averse utility $\log(x)$ her strategy is to spend all the $1$ unit of fund to buy two units of contract $o=1$ whenever $p_{A}=1$, and buy $2$ units of $o=0$ otherwise. With $\epsilon$ probability of losing all her fund, her expected risk-adjusted utility is thus $(1-\epsilon)U(2)+\epsilon U(0)$, which goes to -$\infty$ when $U(x)=0$, regardless how small $\epsilon$ is.
If the decision maker adopts the perfectly calibrated forecaster $B$, she would spend all fund to buy $o=0$ contract when $p_{B}=0$, but split her $1$ unit of fund to buy $2B_{1}$ of $o=1$ contracts and $2B_{0}$ of $o=0$ contracts, so as to maximize her risk-adjusted utility as $\max_{B_{0}+B_{1}=1}2/3U(2B_{1})+1/3U(2B_{0})$. Simple calculation shows that optimal $B_{1}=2/3$, leading to optimal risk-adjust utility $2/3U(4/3)+1/3U(2/3)=\log(\frac{32}{27})/3$, which is strictly larger than her initial utility $\log 1=0$.
Remarks. A few remarks are worthwhile to mention about this example. First, forecaster $p_{A}$ with a small $\epsilon$ will be better for a risk-neural decision maker. However, in this case, while the decision maker will have larger expected utility, she will have to bear the possibility of losing all her $1$ unit of fund, though also enjoy the possibility of doubling her fund to $2$ units. This is often not desirable in real-world applications, such as financial decisions in which risk control is extremely important. In such cases, more well-calibrated forecaters are preferred. Second, our example has extremely negative utility $\infty=U(0)$. This helps us to cleanly highlight the underlying conceptual message, but simple tweak to the utility function can remove that extreme situation yet arrive at the same conclusion.
@@ B.5 A Unified Framework for Utility-Maximizing Betting Strategy
Contracts and market prices. Each binary market resolves to either Yes or No. A share of a Yes contract pays \$1 if the outcome is Yes and \$0 otherwise; the price of this contract is denoted $q$. Symmetrically, a share of a No contract costs 1q1-q and pays \$1 if the outcome is No. These prices are often interpreted as market-implied probabilities (wolfers2006interpreting), though our framework does not rely on this interpretation.
@@ B.7 Bootstrap Confidence Intervals
To assess the uncertainty of the resulting point estimate, we construct a nonparametric bootstrap confidence interval (CI) at level $(1-\alpha)\%$ (diciccio1996bootstrap). Specifically, we form a symmetric interval around the point estimate using bootstrap resampling. A pseudo-code for our implementation is given below in Algorithm 1.
@@
B.8 Robustness to Event-Category Rebalancing
To examine whether our results are sensitive to the distribution of event categories, we perform robustness analyses using rebalanced subsets of the evaluation data. The main dataset is heavily skewed toward Sports events; to assess the impact of this imbalance, we construct alternative evaluation sets in which the proportion of Sports events is reduced while still retaining a mix of the other categories. One subset retains Sports as the largest category but reduces its dominance to roughly 50% (Sports: 0.5, Entertainment: 0.15, Politics: 0.15, Other: 0.2). Another subset ensures that all four categories are equally represented (25% each). All analyses use only submissions on or before 2025–09–10.
Table 5 reports the Brier scores across the original and rebalanced evaluation sets. The relative ordering of the top five models remain unchanged across all distributions, confirming that their comparative performance is robust to event-category composition. As the proportion of Sports events decreases, overall Brier scores improve slightly, suggesting that Sports events are intrinsically more challenging to predict.
Forecaster Original (N=816) Moderate (N=250) Full (N=152)
GPT-5R 0.179 (±0.008)\ ($\pm$ 0.008) 0.146 (±0.013)\ ($\pm$ 0.013) 0.128 (±0.015)\ ($\pm$ 0.015)
Grok 4R 0.186 (±0.009)\ ($\pm$ 0.009) 0.156 (±0.014)\ ($\pm$ 0.014) 0.141 (±0.018)\ ($\pm$ 0.018)
Claude Sonnet 4R (Thinking) 0.190 (±0.009)\ ($\pm$ 0.009) 0.162 (±0.016)\ ($\pm$ 0.016) 0.148 (±0.020)\ ($\pm$ 0.020)
Gemini 2.5 FlashR (Reasoning) 0.195 (±0.009)\ ($\pm$ 0.009) 0.165 (±0.016)\ ($\pm$ 0.016) 0.157 (±0.021)\ ($\pm$ 0.021)
Llama 4 Scout 0.230 (±0.011)\ ($\pm$ 0.011) 0.202 (±0.019)\ ($\pm$ 0.019) 0.210 (±0.025)\ ($\pm$ 0.025)
Market Baseline 0.188 (±0.008)\ ($\pm$ 0.008) 0.164 (±0.010)\ ($\pm$ 0.010) 0.149 (±0.013)\ ($\pm$ 0.013)
Table 5: Brier scores across the original and rebalanced evaluation subsets. Numbers in parentheses denote 95% confidence intervals. The “Moderate” subset contains 50% Sports events, while the “Full” subset has 25% Sports events.
As such, these results provide evidence that the conclusions drawn in the main text regarding LLM performance are robust to variations in the event-category distribution, though we intend to expand the diversity of our events in the future.
@@ C.2 Sharpe Ratio for Average Return
As discussed in § 3.2 and Table 6, Average Return often exhibits wide confidence intervals, reflecting the high variance of event-level payoffs. In most events, models earn nothing or only marginal gains, while in rare cases they achieve outsized returns. To account for this variability, we complement Average Return with the Sharpe ratio (sharpe1998sharpe), a standard metric that normalizes expected returns by their volatility. Formally, for asset returns $R_{a}$, the Sharpe ratio is defined as
@@ C.3 Probability Elicitation Methods
    Verbalized Probability (tian2023just)
@@
    • Prompt emsemble (wightman2023strength): for each market, we average the probabilities elicited from the default and all variation prompts (i.e. (A), (B), (C) above).
@@
    Self-consistency (wang2022self)
@@
    • Weighted (taubenfeld2025confidence): we further supplement each Yes/No with a confidence score. [^14] This confidence reflects uncertainty about the chosen answer, not a direct market probability (e.g., low confidence in Yes signals indecision, not belief in “NO”). Probabilities are formed by confidence-weighted aggregation.
@@ C.6 Evaluating Reasoning using an LLM-as-a-Judge Framework
To directly evaluate LLMs reasoning processes, we employ an LLM-as-a-judge framework (zheng2023judging) to assess the soundness of their reasoning methodologies. Our evaluation proceeds in two stages: (1) eliciting explicit reasoning from the prediction model, and (2) systematically evaluating this reasoning with an independent evaluator.
@@
C.6.1 Validating LLM-as-a-Judge via Human Ratings
To validate the LLM-as-a-judge system and check for alignment bias, we conducted a blinded human adjudication study. We randomly sampled 170 events with LLM-generated reasoning traces and hid both the model identity and LLM judge outputs from human raters. Each reasoning trace was independently scored by humans on a 1-5 scale across the five rubric dimensions (Sources Used, Evidence Extracted, Combination & Weighting, Uncertainties & Counterpoints, and Mapping to Final Probabilities). The rubric provided was identical to that given to the LLMs.
Overall, we find a high degree of concordance between human ratings and LLM-as-a-judge scores. As shown in Table 11, mean absolute differences between human and LLM ratings ranged from 0.37 to 0.44, with an average difference of 0.42, and variances between 0.27 and 0.36. These values are strikingly low given the 1-5 rating scale: all differences are, on average, substantially smaller than 1, meaning that the LLM judge’s score deviates from human judgment by less than half a point. Moreover, across all rubric categories, over 94% of human ratings were within 1 point of the LLM’s score as seen in Fig. 9, underscoring the tight alignment between automated and human evaluation.
Rubric Category Mean Absolute Difference (0-4) (Variance)
Sources Used 0.43 (0.30)
Evidence Extracted 0.42 (0.27)
Combination Weighting 0.43 (0.36)
Uncertainties Counterpoints 0.44 (0.28)
Mapping To Final Probabilities 0.37 (0.29)
Table 11: Human ratings for rubric categories, with variance shown in parentheses. Mean absolute difference is calculated as the average $\lvert\,\text{human score}-\text{LLM score}\,\rvert$ across the 170 events. Differences are on a 0-4 scale.
Figure 9: Agreement matrix showing the percentage of human ratings that lie within exact, moderate ($\pm 1$), and larger ($\pm 2$) thresholds of the LLM-judge ratings for each rubric dimension.
C.6.2 (Lack of) Effect of Scaffolding Prompt on Prediction